solve for x.
log(subscript 3)(x^2+5x)-log(subscript 3)14=0

Answer by edjones(7311)

log[3]((x^2+5x)/14)=0
3^log[3]((x^2+5x)/14)=3^0
(x^2+5x)/14=1
x^2+5x=14
x^2+5x-14=0
(x+7)(x-2)=0
x=-7, x=2
.
Ed